OpenStudy (bmorse):

g(x)=6x^1/3 + 3x^4/3 Identify intervals where the function is increasing/decreasing, concave up/down, critical numbers and inflection points??

OpenStudy (abb0t):

Take the first derivative and the second derivative to find what you seek here. Can you do that ? Then, find the roots.

OpenStudy (bmorse):

I took the derivative and got critical numbers x=-1/2,0 but I'm not sure how to set it up to find where it is increasing and decreasing and so on...

OpenStudy (anonymous):

If you have a graphing calculator you could plug it in and observe that way ...

OpenStudy (bmorse):

there are steps to take and answer using the derivatives...I just don't know exactly how to apply them in this specific case.

OpenStudy (anonymous):

\[g(x)=6x^{\frac{1}{3} }+ 3x^\frac{4}{3}\] \[g'(x)=2x^{-\frac{2}{3}}+4x^{\frac{1}{3}}\] you can do nothing with this until you get rid of the exponents and write what it really is

OpenStudy (anonymous):

\[g'(x)=\frac{2}{\sqrt[3]{x^2}}+4\sqrt[3]{x}\]

OpenStudy (anonymous):

add them up